OpenStudy (anonymous):

Taylor series for ln(2+3x^4)? Find the coefficient on x^44 in the Taylor series for f(x) based at b = 0.

OpenStudy (sirm3d):

first, the taylor series at b=0 of the ln function \[\ln(1+y)=\sum_{n=1}^\infty\frac{(-1)^{n+1}}{n}y^n\]

OpenStudy (sirm3d):

next is to rewrite the given function \[\ln(2+3x^4)=\ln(2\cdot(1+\frac{3x^4}{2}))=\ln2+\ln(1+\frac{3x^4}{2})\]

OpenStudy (sirm3d):

using y=3x^4/2, \[\ln(2+3x^4)=\ln2+\sum_{n=1}^\infty\frac{(-1)^{n+1}}{n}(\frac{3x^4}{2})^n\]

OpenStudy (sirm3d):

the 12th term of the taylor series (n=11) gives you the desired coefficient